Use appropriate algebra and Theorem 7.2.1 to find the given inverse Laplace transform. (Write your answer as a function of t.)

ℒ−1{(1/s2)  − (720/s7)}


1
Expert's answer
2021-04-15T06:56:33-0400

"L^{-1}(\frac{1}{s^2}-\frac{720}{s^7})=L^{-1}(\frac{1}{s^2})-L^{-1}(\frac{6!}{s^7})="


"=t-t^6"


because "L^{-1}(\frac{n!}{s^{n+1}})=t^n"